### v1

# HoneyGPT: Breaking the Trilemma in Terminal Honeypots with Large Language Model

Abstract.
Honeypots, as a strategic cyber-deception mechanism designed to emulate authentic interactions and bait unauthorized entities, continue to struggle with balancing flexibility, interaction depth, and deceptive capability despite their evolution over decades. Often they also lack the capability of proactively adapting to an attacker’s evolving tactics, which restricts the depth of engagement and subsequent information gathering. Under this context, the emergent capabilities of large language models, in tandem with pioneering prompt-based engineering techniques, offer a transformative shift in the design and deployment of honeypot technologies. In this paper, we introduce HoneyGPT, a pioneering honeypot architecture based on ChatGPT, heralding a new era of intelligent honeypot solutions characterized by their cost-effectiveness, high adaptability, and enhanced interactivity, coupled with a predisposition for proactive attacker engagement. Furthermore, we present a structured prompt engineering framework that augments long-term interaction memory and robust security analytics. This framework, integrating thought of chain tactics attuned to honeypot contexts, enhances interactivity and deception, deepens security analytics, and ensures sustained engagement.
The evaluation of HoneyGPT includes two parts: a baseline comparison based on a collected dataset and a field evaluation in real scenarios for four weeks. The baseline comparison demonstrates HoneyGPT’s remarkable ability to strike a balance among flexibility, interaction depth, and deceptive capability. The field evaluation further validates HoneyGPT’s efficacy, showing its marked superiority in enticing attackers into more profound interactive engagements and capturing a wider array of novel attack vectors in comparison to existing honeypot technologies.
1. Introduction
Honeypots have been widely used as effective cybersecurity tools for detecting, enticing, and understanding malicious activities on the Internet  (López-Morales et al., 2020; Kim et al., 2022; Sasaki et al., 2022; Anagnostakis et al., 2005; Wahab et al., 2019; Rrushi, 2018; Torabi et al., 2020; Zhan et al., 2013; Wang et al., 2023; Huang and Zhu, 2021). These traps are designed to lure attackers, allowing security teams to gather their information and monitor their behaviors for threat analysis, as well as
deflect ongoing attacks.
Terminal honeypots are a specialized type of honeypots for emulating real terminal systems. They are typically employed to attract hackers or malicious software targeting terminal systems. Terminal honeypots range from low-interaction variants that simulate a minimal set of protocols or service information  (Niels, 2004), to medium-interaction honeypots that replicate a broader array of system functionalities at a higher deployment and maintenance cost, and high-interaction honeypots that leverage sophisticated virtualization technologies, such as VMWare to create convincing environments of terminal systems or even deploy actual systems for maximum authenticity.
However, existing terminal honeypots face a trilemma problem related to flexibility, interaction level, and deceptive capability, stemming from the limitations of programmatic approaches and the fixed nature of system environments. In the trilemma, existing terminal honeypot designs struggle to achieve optimal performance across all three dimensions simultaneously, forcing defenders to compromise and select imperfect solutions.
It is very challenging to address the trilemma in practice. The primary methods for running honeypots are based on either programmatic simulations or real operating systems. On one hand, the programmatic simulation approach is limited by development costs, making it difficult to balance scalability and interaction depth. On the other hand, honeypots based on real operating systems suffer from rigidity, due to their inherent configurations that are short of flexibility. Moreover, both deployment methods lack intelligence during interactions, unable to provide responses that align with attackers’ intentions to entice further engagement. This significantly limits the honeypot’s deception capabilities. However, there has been a growing demand for honeypots that are more dynamic, intelligent, and cost-effective.
The development of Large Language Models (LLMs) (Kaplan et al., 2020; Wei et al., 2022a; Fu et al., 2022; Feng et al., 2024; Yang et al., 2022b) appears to offer a potential solution to these problems. By providing attack commands as queries to the large language model, it can generate responses tailored to the attackers’ interests based on the provided prompt. This approach breaks the Trilemma of traditional honeypots, as a single LLM can simultaneously simulate multiple high-interaction honeypots with different configurations. Furthermore, by designing prompts to guide the LLM, it is possible to better fulfill attackers’ objectives and entice them into deeper levels of interaction.
However, designing honeypots based on LLMs still faces three major challenges. A prominent issue is the absence of a standardized approach to honeypot prompt design. To address this problem, this paper establishes a universal specification for honeypot prompt keywords. Another challenge arises from the native thinking constraints of LLMs, which cannot handle complex attack combinations. To address this challenge, we employ a Chain of Thought (CoT) strategy that allows the language model to answer the command’s impact on the operating system during each interaction. When posing the next question, we integrate the consolidated changes in the operating system and the user command to enhance the large language model’s understanding of the task. The third challenge lies in that the performance of LLMs in prolonged sessions is hampered by their context length limitations and inherent forgetting mechanisms. To manage this issue, we employ the CoT approach to assess the significance of each attacker command and its impact on the terminal system. This assessment allows us to strategically prune the session history, ensuring that only the most essential contextual information is retained within the confines of the context length, thereby optimizing the honeypot’s effectiveness in extended interactions. The major contributions of this paper are summarized as follows.
1. We propose HoneyGPT, an intelligent terminal honeypot that addresses the dimensions of flexibility, interaction level, and deceptive capability, enabling extended interactions while overcoming the Trilemma.
2. We propose a Chain of Thought strategy for terminal honeypot task scenarios, which enhances the task-solving capabilities of native LLMs and overcomes the inherent context length limitations. The introduction of CoT is essential to advancing honeypot technology and improving network security defense capabilities.
3. We conduct a baseline evaluation of HoneyGPT in comparison with traditional honeypots, utilizing open-source attack data captured by Cowrie. The results indicate that HoneyGPT outperforms conventional honeypots with respect to its composite capabilities in flexibility, level of interaction, and deception efficacy.
4. We deploy HoneyGPT and Cowire, which is a programmatic honeypot, on the Internet for four weeks and observe that HoneyGPT captures more attack behaviors and longer interactions than Cowrie.
The rest of this paper is structured as follows. Section 2 provides background on LLMs, prompt engineering, and terminal honeypots. Section 3 describes the trilemma of flexibility, interaction, and deception in existing terminal honeypots. Section 4 presents our HoneyGPT solution, detailing its framework, concept, strategy, prompt management, and configuration. Section 5 validates the efficacy of HoneyGPT, providing baseline comparisons on deception, interaction, and flexibility, as well as field evaluations to measure its effectiveness in engaging attackers and capturing novel attack vectors. Section 6 outlines the limitations and potential future work, and finally, Section 7 concludes the paper.
2. Background
In this section, we provide the background of Large Language Models, prompt engineering, and terminal honeypots.
2.1. Large Language Models
Large Language Models (LLMs) are a subset of pre-trained language models (PLMs), characterized by their voluminous parameter sets, reaching or even exceeding hundreds of billions, as defined by Kaplan et al (Kaplan et al., 2020). Since OpenAI introduced ChatGPT  (openai, [n. d.]), LLMs have been widely used across a spectrum of disciplines  (Fu et al., 2022; Brown et al., 2020), particularly in the field of cybersecurity  (Meng et al., 2024; Hu et al., 2024; He and Vechev, 2023; Pearce et al., 2023, 2022; Vishwamitra et al., 2024). Compared to previous models, LLMs offer two main advantages in interdisciplinary applications:
Firstly, LLMs display astonishing emergent abilities as identified by Wei et al (Wei et al., 2022a), which are competencies absent in smaller-scale models but manifest in larger configurations. Such models experience a quantum leap in performance relative to their diminutive counterparts. Prominent among these emergent abilities are:
  • In-context learning: LLMs can, upon receiving natural language instructions or several task demonstrations, generate accurate outputs for test cases by continuing the input text sequence, circumventing the need for further training or gradient adjustments  (Brown et al., 2020).
  • Instruction following: LLMs possess the ability to comprehend and execute natural language instructions with minimal or even zero examples, adapting to new tasks (Victor et al., 2022). By fine-tuning tasks articulated through instructional prompts, LLMs have demonstrated proficiency in accurately handling tasks they have never seen before (Ouyang et al., 2022).
  • Step-by-step reasoning: Distinct from conventional PLMs, LLMs are capable of deconstructing complex tasks into a sequence of discrete subtasks by employing prompt-based strategies such as chain-of-thought. It is conjectured that this capacity may be derived from their training on code  (Wei et al., 2022b; Fu et al., 2022).
Secondly, LLMs transform the traditional employment of AI algorithm development and utilization, significantly diminishing the technical threshold for cross-disciplinary actors to harness such algorithms. In contrast to previous models, users engage with LLMs via a prompt interface (e.g., the GPT-3-turbo API), by formulating natural language prompts or directives to modulate the model’s behavior to yield anticipated results. This empowers individuals who lack a deep understanding of model training paradigms but are conversant with prompt engineering to employ LLMs for task facilitation.
2.2. Prompt Engineering
Prompt engineering is the process of crafting prompts to better guide large language models to understand and resolve tasks. Here we briefly introduce prompt engineering from its core facets: components, design principles, and strategic approaches.
Components
The standard components of LLM prompts typically include:
  • Task Description: Instructions that the LLM is expected to adhere to, expressed in natural language.
  • Input Data: The requisite data for a task, presented in natural language.
  • Contextual Information: The contextual or background information that can aid in the execution of a task.
  • Prompt Style: The formulation of the prompt, which is tailored to optimize the model’s responses, such as through role-playing or incremental reasoning.
Design Principles
The development of effective LLM prompts is predicated on:
  • Clarity: Ensure instructions are explicit and communication is clear through symbolic delimiters, structured outputs, conditional logic, and ”Few-shot” examples.
  • Deliberation: Allow LLMs time for thought, structuring prompts for sequential processing and reasoning with ”chain of thought” techniques. Allow LLMs time to reason by structuring prompts for sequential information processing using ”chain of thought” techniques.
Prompting Strategies
The ”Chain of Thought” (CoT)  (Wei et al., 2022b; Feng et al., 2024; Yang et al., 2022b; Ling et al., 2024; Lu et al., 2022; Yang et al., 2022a; Zelikman et al., 2022) paradigm has been shown to significantly enhance LLM performance on tasks that require complex reasoning by incorporating intermediary reasoning steps within the prompts. The zero-shot CoT, as introduced in foundational research, employs prompts such as ”Let’s think step by step” to shepherd LLMs to a logical conclusion, thus highlighting the significance of scale in emergent capabilities.
2.3. Terminal Honeypots
Honeypots are security tools designed as decoys to attract and monitor unauthorized or malicious attackers. Based on their implementation approaches, terminal honeypots can be categorized into three primary types: emulated honeypots, which are based on programmed simulations; real system honeypots, which utilize real systems; and hybrid honeypots, which combine the features of the first two types.
Emulated Honeypots
Emulated honeypots are software constructs that simulate network services and devices to attract cyber adversaries. They are developed using conventional programming methodologies, as exemplified by Honeyd  (Niels, 2004), Dionaea (phibos, [n. d.]), and Cowrie (OOSTERHOF, [n. d.]). These honeypots utilize rule-based matching and logical controls to mimic desired targets. Despite their effectiveness, they are constrained by the limitations of development and operational costs, preventing a full emulation of an operating system’s intricate behaviors. The interactivity spectrum of these honeypots spans from minimal to moderate: Honeyd and Dionaea deliver low-interaction platforms rich in configuration possibilities, while Cowrie steps up to offer a medium-interaction milieu. However, this comes at the cost of diminished adaptability and increased operational costs. Such frameworks have laid the groundwork for a range of specialized offshoots (López-Morales et al., 2020; Alexander and Richard, 2018).
Real-system Honeypots
Real-system honeypots, by contrast, utilize genuine, unmodified system environments, offering a more authentic simulation of user interactions, file system activities, and operational behaviors  (Wang et al., 2022; Anagnostakis et al., 2005; Wahab et al., 2019). They provide a high level of interaction, creating a convincing facade of a real system, which significantly increases the probability of capturing complex attack strategies that simpler honeypot architectures may miss. However, the inherent rigidity of their physical environments (Alexander and Richard, 2018) requires individualized configuration for each new system variant, which incurs substantial deployment and maintenance costs.
This authenticity offers potential attackers a highly interactive environment, fostering a convincing deception that aids in uncovering intricate attack vectors that might elude detection by more rudimentary honeypots. However, a challenge arises from the security protocols native to real operating systems, such as stringent permission controls, which are not inherently conducive to fulfilling malicious requests. These protocols can inadvertently deter threat actors from engaging in more profound levels of system exploitation, thus potentially curtailing the honeypot’s effectiveness as a lure. Additionally, the deployment and maintenance of real-system honeypots are encumbered by the need for individual configurations for each system variant, leading to significant resource expenditures.
Hybrid Honeypots
In an effort to combine the strengths of both emulated and real system honeypots, a scalable hybrid honeypot framework has been proposed (Xu et al., 2021; You et al., 2021). This framework employs programmed methods for simulating low-level interactions, such as network protocols, while delegating high-level interactions, such as file executions, to a real file system. Although the hybrid model ostensibly enhances flexibility and interaction depth, it is not devoid of limitations. The framework may experience discordance between the simulated protocols and the actual file system responses, leading to a detectable discrepancy that astute attackers could potentially exploit. Furthermore, the rigidity in configuring a physical environment restricts adaptability and diminishes the diversity of supported environmental configurations, which could impair its capacity to effectively deceive sophisticated attackers.
3. Trilemma of Existing Terminal Honeypots
The trilemma in honeypot design reflects the intricate balancing act required to optimize the three key attributes of effective honeypot systems: interaction depth, flexibility, and deception ability. Honeypots from their inception have grappled with this trilemma, consistently facing trade-offs that impede their overall efficacy.
3.1. Limited Flexibility
The narrow concept of flexibility within honeypot design is indicative of the system’s adaptability in authentically replicating a variety of systems and services. Such adaptability is pivotal for honeypots, as it underpins their ability to create authentic and persuasive simulations of various operational environments. Yet, the quest for such flexibility is one of the most formidable challenges in honeypot development, with the central issue being the support for a wide range of operating system versions and nuanced system configurations.
Contemporary honeypots often struggle to provide sufficient flexibility, particularly regarding the support of diverse operating system versions and security configurations. The pursuit of increased flexibility can compromise the depth of interaction and the honeypot’s deception ability.
Code-based emulated honeypots exemplify this issue, as they require extensive resources for the creation and maintenance of simulations that encompass numerous operating system versions and environmental configurations. Such complexity leads to substantial development and operational costs, and no code-based honeypot has yet managed to seamlessly transition across varied terminal environments while maintaining high interaction fidelity. Real-system honeypots offer detailed interactions but lack the adaptability of their emulated counterparts, making them difficult to reconfigure in response to evolving threats. They also impose significant management overhead, due to the requirement of maintaining real operating system environments.
Moreover, traditional honeypot offerings remain constrained to single-dimensional deception techniques. As to leveraging the data they capture, there is a conspicuous lack of flexibility in integrating security analysis, necessitating distinct systems dedicated to data processing and analysis.
The core challenge, therefore, is to architect honeypots that strike an optimal balance between delivering authentic, complex interactions and possessing the agility to counter novel attacks. Achieving this equilibrium should not exacerbate costs or inflate management burdens, crafting a solution that is as economical as it is effective.
3.2. Limited interaction
Due to limitations in implementation methods and development costs, code-based terminal honeypots are restricted to simple rule matching and lack the capability to execute complex commands as a real operating system. This limitation significantly reduces the level of interaction these honeypots can manage, making them inadequate in responding to complex commands issued by attackers. For instance, Cowrie, a commonly used middle-interactive honeypot, is unable to handle complex commands that involve concatenation with pipe symbols (—). The inability to process complex operations can lead to honeypot instability or crashes, thereby failing to further engage attackers. Enhancing the interaction of honeypots is crucial for increasing their reliability and effectiveness.
3.3. Limited Deception
A key factor contributing to limited deception is the lack of dynamic and adaptive behavior in existing honeypot systems. Code-based terminal honeypots typically rely on static configurations and predefined matching mechanisms, while honeypots based on real operating systems often perform limited and monotonous simulations of the operating environment due to cost constraints. Experienced attackers can easily identify these singular and rigid environment configurations. The lack of variability reduces the effectiveness of deception because attackers can quickly recognize and bypass the simulated environment of the honeypot. Additionally, code-based terminal honeypots lack authentic user interactions and the realism of contextual information, further limiting their ability to convincingly simulate real systems. When attackers perform complex operations such as file execution or process communication, code-based honeypots struggle to replicate authentic user interactions or provide the depth of a real system.
In addition, in terms of enticing attackers for deeper-level attacks, current honeypots often fall short. Whether it is a code-based honeypot like Cowrie or a honeypot based on real systems, the limited configurations and restricted permissions provided to attackers often fail to meet the intentions of attackers. Catering actions such as modifying user passwords, terminating competing processes, or viewing graphics card information, which could effectively boost an attacker’s motivation, are often overlooked by existing honeypots. Especially in a real system environment, the operating system’s characteristics, such as permission control mechanisms and limited configurations (such as hardware devices and account password configurations), have limitations in fulfilling the attacker’s attack intentions.
4. HoneyGPT
To address the trilemma of existing honeypots, we introduce HoneyGPT, an advanced honeypot framework based on ChatGPT. This framework transforms the conventional ”request-response” message interaction (based on terminal protocols) into a ”question-answer” text interaction (based on ChatGPT API).
As shown in Figure  1, the framework consists of three main components: the Terminal Protocol Proxy, the Prompt Manager, and ChatGPT. The Terminal Protocol Proxy is responsible for fundamental protocol tasks, such as establishing connections, fingerprint emulation, message encapsulation, and parsing. We utilize the protocol layer of the Cowrie honeypot to develop the Terminal Protocol Proxy, which supports protocols such as SSH and Telnet. The Prompt Manager is tasked with constructing prompts based on commands sent by attackers and interacting with ChatGPT. ChatGPT generates responses based on the provided prompts. After receiving a response from ChatGPT, the Prompt Manager extracts the terminal output and forwards it to the Terminal Protocol Proxy, which then encapsulates it into messages and returns them to attackers. Since the Terminal Protocol Proxy repurposes the work of the Cowrie honeypot and is not a primary innovation, we will predominantly elaborate on the design of the Prompt Manager in the following section.
Figure 1. HoneyGPT Framework
Figure 2. HoneyGPT Architecture
4.1. Concept and Strategy
We first articulate the fundamental concepts and methodological strategies that are integral to the design of the Prompt Manager.
4.1.1. Concept of HoneyGPT
We define the interaction history as a sequence of N question-answer pairs:
(1) $\{(Q_{1},A_{1}),(Q_{2},A_{2}),\ldots,(Q_{N},A_{N})\}$
To obtain the response $A_{i}$ from the $i$-th interaction, we constructed complex prompts. The operational logic of HoneyGPT during the $i$-th interaction is showcased in Figure 2. The $i$-th HoneyGPT interaction is formally defined by Equation  2.
(2) $(A_{i},C_{i},F_{i})=ChatGPT(P,S,H_{i},SR_{i},Q_{i})$
The components $P$, $S$, $H_{i}$, $SR_{i}$, and $Q_{i}$ together constitute the prompt for the $i$-th interaction. Here $P$ and $S$ represent the static configuration of the prompt, remaining unchanged throughout the interactions. $Q_{i}$ is the command sent by the attacker during the $i$-th interaction. $H_{i}$ and $SR_{i}$ are dynamic aspects of the prompt, updated after each interaction by the Prompt Manager M. The Prompt Manager synthesizes these variables into the new prompt for subsequent dialogue sessions, as defined by:
(3) $M(P,S,Q_{i},SR_{i},C_{i},F_{i},A_{i},w)=(P,S,H_{i+1},SR_{i+1},Q_{i+1})$
The following are definitions for each variable:
Honeypot Principle $P$ defines the strategies and guidelines used to create a honeypot.
Honeypot Setting $S$ refers to the initial configuration set within the prompt for the Honeypot.
History of Interaction $H_{i}$$\{(Q_{1},A_{1}),(Q_{2},A_{2}),\ldots,(Q_{i-1},A_{i-1})\}$ provides ChatGPT with the essential context required to support its capacity for extended interaction memory. To ensure that the length of $H_{i}$ remains within the context length limits of ChatGPT, the Prompt Manager $M$ implements a pruning algorithm to regulate the size of the historical context. This algorithm judiciously trims components of $H_{i}$ that have less significance, striking a balance between retaining pertinent context and maintaining the model’s operational efficiency within its architectural constraints.
System State Register $SR_{i}$ refers to a collection of system states communicated to ChatGPT at the $i$-th interaction: $\{(C_{1}),(C_{2}),\ldots,\\ (C_{i-1})\},$ where each state $C_{i}$ corresponds to the system’s changes after each prior interaction, paralleling the order of $H_{i}$. $SR_{i}$ is the manifestation of a ’chain-of-thought’ strategy, which HoneyGPT leverages to enhance its capacity for generating the subsequent response $A_{i}$.
Attacker Query $Q_{i}$ in the context of HoneyGPT, refers to the command issued by the attacker to the honeypot terminal system during the $i$-th interaction.
Honeypot Answer $A_{i}$ in the context of HoneyGPT, denotes the command execution echo returned by the honeypot terminal to the attacker during the $i$-th interaction.
System State Evaluation $E_{i}$ is an assessment provided by ChatGPT regarding the system’s state post-interaction, consisting of two parts: Impact Factor $F_{i}$, which values the impact of $Q_{i}$ on the system, and $C_{i}$, which denotes the changes to the terminal system environment after $Q_{i}$ execution.
Weaken Factor $w$ is the constant used in the pruning algorithm, which accounts for the timing of interactions, preferentially trimming older and less relevant exchanges to maintain a recent, pertinent dialogue history.
4.1.2. Strategy
Figure 3. Comparison of the Effects of Integrating CoT
The design of HoneyGPT incorporates a Chain of Thought (CoT) strategy to address the intrinsic limitations of ChatGPT in processing complex, extended dialogues and to circumvent the constraints of the prompt context length. The essence of CoT lies in compelling the LLM to decompose a multifaceted problem into a sequence of sub-problems, tackling them incrementally to enhance the LLM’s ability to comprehend and address intricate issues.
Complex Extended Dialogue Tasks: In realistic adversarial scenarios, attackers often execute a series of malicious activities through multiple commands. As shown in Figure 3, attackers issue the ”write-elevate-execute” attack sequence. While native ChatGPT can effectively handle individual commands, it struggles with complex, interdependent, extended dialogue tasks that require memorization and coherence. By contrast, HoneyGPT can significantly increase the capability to handle such complex command sequences by incorporating the CoT strategy.
Within HoneyGPT, we simplify the complexity of extended dialogues into two steps: first, assessing the impact of previous commands on the operating system, and then determining the system’s response to new adversarial commands based on the evaluated impact. To ensure the continuity and automation of the attack interaction process, the task of assessing the impact of attack commands on the operating system has been delegated to ChatGPT. Thus, during each interaction, the prompt must provide ChatGPT with details of the changes that have occurred within the system up to that point( $SR_{i}$ ), as well as the present attack command( $Q_{i}$ ). ChatGPT is then tasked with analyzing the expected output from the terminal ( $A_{i}$ ) and determining the result of System State Evaluation ( $C_{i}$, $F_{i}$ ). Following the conclusion of an interaction, the Prompt Manager integrates $C_{i}$ into $SR_{i}$, respectively, setting the stage for the subsequent interaction.
Prompt Pruning Strategy: After each interaction, the Prompt Manager appends content to prompt, respectively, which results in an increasingly lengthy prompt for subsequent interactions. However, due to limitations imposed by computational resources, storage capacities, and model characteristics, LLMs have a restricted context length for prompts. Should the prompt’s token count exceed this threshold, ChatGPT will be unable to function. To address this issue, we introduce a strategy known as Prompt Pruning: if the prompt’s context length surpasses the limit, we eliminate the least impactful entries from the dynamic portion, $H_{i}$, with respect to the current operating system. Although $SR_{i}$ also grows dynamically, its length is significantly shorter than the context length limit, and therefore, it does not require pruning. Since $H_{i}$ represents the collection of historical honeypot interactions, it suffices to consider both the temporal sequence and the magnitude of impact to identify and remove the least pertinent interaction records for the current exchange. The impact magnitude is determined by ChatGPT during each interaction, for which we have provided a set of grading criteria that ChatGPT uses to generate $F_{i}$, as listed in Table 1.
4.2. Prompt Manager $M$
The operations of the Prompt Manager are categorized into three main processes: prompt construction, prompt pruning, and content updating.
Table 1. Rules of the Assignment of Values to $F_{i}$
Value Condition
0 Read file, display system information
1 file created, install tool
2 modify files/dir, change working directory, change shell
3 Start/stop service, download file, Elevate privilige
4 Impact services, delete files, password changed
4.2.1. Prompt Construction
The construction process takes place before each interaction session. When the attacker issues a command to HoneyGPT, the Prompt Manager formulates three critical questions for ChatGPT to answer, based on the Honeypot Principle, Honeypot Setting, Attacker Query, System State Register, and History of Interaction:
  1. (1)
    What is the terminal’s response ($A_{i}$) ?
  2. (2)
    How does the system state change ($C_{i}$) ?
  3. (3)
    Assign a numerical value to the command’s aggressiveness ($F_{i}$).
Question 2 underpins the CoT strategy, while Question 3 is intended for prompt length pruning and will be discussed later. Upon obtaining responses Ai, Ci, and Fi, HoneyGPT relays Ai back to the attacker, while Ci and Fi are utilized by the Prompt Manager to update the Hi and SRi in prompt for the subsequent interaction.
4.2.2. Prompt Pruning
Pruning is activated when the length of the prompt exceeds the operational limits of commercial LLMs. The primary objective is to trim interactions of less importance from the History of Interaction ($H_{i}$). Since the entries regarding system state changes are typically concise, $SR_{i}$ are exempt from this curtailment. The success of pruning depends on accurately singling out the least significant interactions from $H_{i}$. We have graded the impact of adversarial commands on the current system by using a scale from 0 to 4. The criteria for assigning values to the Impact Factor ($F_{i}$) are detailed in Table 1. LLMs will generate responses by referencing the table, ensuring that the pruning process is executed in a systematic and uniform manner.
In addition, the design contemplates the temporal dimension of command relevance, positing that earlier exchanges bear less pertinence to the present interaction. The Weaken Factor $w$ is introduced to quantify this temporal effect. With each interaction’s conclusion, the Impact Factor $F_{i}$ for each entry in $H_{i}$ is attenuated by multiplication with $w$, effectuating a time-based diminution of $F_{i}$. During the pruning process, entries with the lowest $F_{i}$ values are prioritized for removal to manage and preserve the prompt token limit efficiently.
Regarding the determination of the Weaken Factor $w$, both the length of the dialogue and the relevance of the content are considered. Observations from cybersecurity practices indicate that a sequence of attacks with strong correlation typically progresses through a ”write-elevate-execute” triad. Therefore, it is advisable to ensure that a newly executed command has a higher significance than a high-privilege command that occurred three steps prior, signified by $F_{i+3}\cdot w^{3}<F_{i}$. Additionally, the value of $w$ should not be too diminutive in order to adequately reflect the effect of time on the relevance of past commands. The process of temporal decay is implemented during Content Updating, where $F_{i}$ of each element is multiplied by the Weaken Factor ($w$).
4.2.3. Content Updating
The updating process is conducted at the end of each interaction session, aiming to renew the dynamic cache contents, specifically the System State Register $SR_{i}$ and the History of Interaction $H_{i}$. This does not include static elements such as constants S, P, and the adversary’s instructions $A_{i}$. Upon the completion of the $i$-th interaction, the Prompt Manager incorporates the query-response pair $(Q_{i},A_{i})$ into $H_{i}$, and the context $C_{i}$ into $SR_{i}$. The $SR_{i}$ supports the CoT strategy, while $H_{i}$ serves as contextual information for enriching ChatGPT’s understanding of the interaction history. Additionally, the updating includes the application of a temporal decay to the Impact Factor $F_{i}$ to support the pruning mechanism.
4.3. Configuration of HoneyGPT
Here we introduce the static components involved in configuring HoneyGPT, namely the System Principles ($P$) and Honeypot Settings ($S$), which remain unchanged once a honeypot is initialized. System Principles provide behavioral guidelines for simulating system activities, covering aspects such as the role of HoneyGPT, time sensitivity, input/output formats, and few-shot learning. These principles enable HoneyGPT to effectively attract and analyze potential attack behaviors by maintaining a credible system illusion. The Honeypot Settings involve specific configurations like hardware specifications and software environments, which are essential for simulating realistic targets that are attractive to attackers. These settings are designed to enhance the honeypot’s attractiveness and precision in capturing malicious exploitation such as cryptocurrency mining. Once set, both components provide a stable platform for HoneyGPT to operate effectively within its designed parameters. A detailed description of these two sets is given in Appendix A.
5. Evaluation
Our evaluation of HoneyGPT includes two parts: the baseline evaluation and the field evaluation. The baseline evaluation compares the capabilities in deception, interaction level, and flexibility between HoneyGPT and traditional honeypots. Specifically, to evaluate the honeypots’ capability in deception and interaction level under consistent attack scenarios, the baseline evaluation utilizes the attack dataset collected by Cowrie. The field evaluation involves deploying HoneyGPT and Cowrie simultaneously on the Internet for a period of four weeks to assess whether HoneyGPT can influence attackers’ attack strategies and capture a wider range of attack types and interaction lengths.
5.1. Baseline Comparison Set
The comparison in the baseline evaluation is from three aspects: deception, interaction level, and flexibility. The comparisons of deception and interaction level are performed in a quantitative manner, by utilizing the same attack dataset to ensure the same attack scenarios, while the flexibility comparison is performed in a qualitative manner, through individual experimental verification for each indicator.
Table 2 shows the experimental setup for the baseline comparison. The comparison of deception is conducted among HoneyGPT based on GPT-3.5 turbo interface, HoneyGPT based on GPT-4 turbo interface, Cowrie, and a real system. The interaction level comparison is limited to the comparisons among HoneyGPT based on GPT-3.5 turbo interface, HoneyGPT based on GPT-4 turbo interface, and Cowrie, because the real system exhibits the best interaction level and there is no need to compare. The flexibility comparison is conducted among Honeyd, a low-interaction honeypot, along with HoneyGPT, Cowrie, and the real system.
To assess the differences in deception and interaction level between HoneyGPT and traditional honeypots under the same attack scenarios, we use the same attack dataset to replay traffic, simulating identical attack scenarios. This dataset compiles attack data previously captured by Cowrie, a mainstream, code-based, medium-interaction open-source honeypot system  (Priya and Chakkaravarthy, 2023; Başer et al., 2021). As shown in Table 3, these datasets include a substantial volume of attack commands and valid attack sessions. A sequence of actions performed by an attacker post-successful login is considered as a valid attack session. We perform data cleaning and aggregation on these datasets, resulting in 160 attack sessions and 1,489 attack commands, which are used for baseline comparison.
Table 2. Baseline Comparison Set
Capability Type Subjects Method Research Type
Deception HoneyGPT (GPT-3.5 turbo), HoneyGPT (GPT-4), Cowrie, Real System Replay attack dataset Quantitative Research
Interaction HoneyGPT (GPT-3.5 turbo), HoneyGPT (GPT-4), Cowrie Replay attack dataset Quantitative Research
Flexibility HoneyGPT, Cowrie, Honeyd, Real System Verification Individually Qualitative Research
Table 3. Attack Data Source
Source Capture Method Time Span Number of Attacks Effective Attack Sessions
 (Başer et al., 2021) Cowrie captured 2021/04/06-2021/06/11 5099153 2479
 (Priya and Chakkaravarthy, 2023) Cowrie captured 2022/06/09-2022/10/29 1261689 7374
5.2. Baseline Comparison of Deception
Here we assess the deception capabilities of Cowrie, HoneyGPT based on the GPT-3.5 Turbo API, HoneyGPT based on the GPT-4 API, and a real operating system based on virtual machines.
5.2.1. Evaluation Metrics
We assess the deception of honeypots from two aspects:
  1. (1)
    Whether the execution of commands succeeds or not. Attack behaviors typically exhibit temporal dependencies and control logic, where the execution of commands relies on the desired effects of previous attack commands, such as successful execution or system information that is of interest to attackers. Thus, a higher execution success rate indicates the honeypot’s capability to induce attackers into deeper interactions and capture more valuable information.
  2. (2)
    Whether the outputs of honeypot comply with the OS logic or not. Whether the honeypot’s outputs align with the logic of the operating system directly affects the effectiveness of its deception. Deviations from the expected behaviors of an operating system, such as inconsistent shell prompts or command execution results, may raise suspicion by attackers and expose the honeypot. A higher level of OS logic compliance indicates a stronger deceptive capability of the honeypot.
Based on these two aspects, we then define four distinct categories to classify and compare the deceptive performance of the honeypot system’s responses.
  • SALC (Successful Attack with Logic Compliance): Successful attacks that conform to the logic of the operating system.
  • SALNLC (Successful Attack without Logic Compliance): Successful attacks that do not adhere to the logic of the operating system.
  • FALC (Failed Attack with Logic Compliance): Failed attacks that conform to the logic of the operating system.
  • FALNLC (Failed Attack without Logic Compliance): Failed attacks that do not adhere to the logic of the operating system.
Figure 4. Distribution of Deception Categories Across Different Honeypots
Figure 5. Performance Comparison
With these defined categories, we can now introduce four key evaluation metrics to assess the honeypot system’s deceptive capabilities:
  • Accuracy: It measures the system’s ability to generate outputs consistent with the logic of a real operating system. It is calculated by dividing the number of successful attacks that conform to the operating system’s logic (SALC) by the total number of successful attacks (SALC + SALNLC). Improving accuracy enables more precise simulation of operating system behaviors, lowers attacker suspicion, and enhances information gathering and threat awareness capabilities.
  • Temptation: It evaluates the effectiveness of the system in attracting attackers’ engagement. It is calculated by dividing the number of successful attacks that conform to the operating system’s logic (SALC) by the total number of attacks with operating system logic (SALC + FALC). A higher level of temptation indicates a more successful lure for attackers and greater information collection.
  • Attack Success Rate: It measures the overall success rate of executing deceptive commands. It is calculated by dividing the total number of successful attacks by the total number of attempted attacks. A higher success rate implies more successful interactions and better information collection.
  • OS Logic compliance: It evaluates the system’s capability to generate outputs consistent with the logic of a real operating system. It is calculated by dividing the number of attacks that conform to the operating system’s logic (SALC + FALC) by the total number of attempted attacks. Improving OS logic compliance allows for the creation of more accurate simulations of operating system behaviors, lowering attackers’ suspicions.
5.2.2. Performance Comparison
Figures 4 and 5 illustrate the performance comparison among Cowrie, a real operating system, and HoneyGPT based on GPT-3.5 and GPT-4.
In terms of attack success rate and temptation, HoneyGPT based on GPT-3.5 and GPT-4 outperforms both Cowrie and the real operating system. The inherent characteristics of LLMs enable HoneyGPT to better cater to attackers’ intentions, enticing them to further engage in attacks, which is a significant advantage of HoneyGPT. Furthermore, increasing the model parameter size leads to improvements in both metrics. Regarding OS logic compliance and accuracy, HoneyGPT based on GPT-3.5 and GPT-4 also demonstrates superior performance than Cowrie and approaches the level of the real operating system. Additionally, increasing the model parameter size leads to enhancements in both metrics. Note that with the increase in model parameter size, there is a notable improvement in the proportion of successful execution of deceptive commands that comply with the operating system logic (SALC). This indicates that as the model evolves, the precision of HoneyGPT in catering to attackers’ intentions also increases. In summary, the performance comparison highlights the advantages of HoneyGPT based on GPT-3.5 and GPT-4 over Cowrie and the real operating system. HoneyGPT demonstrates superior performance in terms of accuracy, temptation, attack success rate, and compliance with operating system logic. Moreover, increasing the model parameter size can further improve the performance in these metrics.
5.3. Baseline Comparison of Interaction
Unlike the deception assessment, we consider a honeypot providing interactions as long as it can respond without crashing, even if the honeypot’s feedback suggests that the service does not exist or there is insufficient permission. We assess the interaction level of Cowrie, HoneyGPT based on the GPT-3.5 Turbo API, and HoneyGPT based on the GPT-4 API. The baseline comparison of interaction level can be divided into two parts. First, we compare and analyze the degree of interaction support that HoneyGPT and Cowrie provide in response to diverse attack requests. Second, based on the ATT&CK matrix, we derive the response rates of HoneyGPT to different attack techniques under varied interface supports and analyze the reasons for non-responses to each attack technique.
5.3.1. Performance Comparison
Table 4. Evaluation of Interaction Level
Metrics Cowrie HoneyGPT (GPT-3.5-turbo) HoneyGPT (GPT-4)
Full Session Response Rate 81.88% 93.75% 99.38%
Command Response Rate 96.98% 99.13% 99.93%
Mean Session Length Percentage 57.28% 71.73% 99.93%
Mean Interaction Degree Percentage 83.24% 94.45% 99.91%
To assess the interaction level of a honeypot, we introduce four key metrics:
  • Complete Session Response Rate: The ratio of sessions that are completely and successfully responded to requests throughout their duration.
  • Command Response Rate: The percentage of individual commands within sessions that receive responses.
  • Mean Session Length Percentage: Defined as the ratio of the average executed session length to the overall average session length.
  • Mean Interaction Degree Percentage: Defined as the average of the proportions of successful responses within individual sessions.
Table 4 presents a detailed comparison of the interaction levels for each honeypot. The results clearly demonstrate the superior performance of HoneyGPT, especially the version based on the GPT-4 API, achieving nearly perfect interaction performance across all categories. This indicates the superiority of HoneyGPT’s capability to maintain interactive sessions and support command interactions.
5.3.2. Unresponsiveness Analysis and Error Origins
We assess the interactive performance of HoneyGPT based on the GPT-3.5 Turbo API and that of HoneyGPT based on the GPT-4 API in responding to various attack vectors, analyzing the types and frequencies of errors they encounter with each type of attack.
Our analysis employs the ”Techniques” component of the ATT&CK framework developed by MITRE to classify the test data (Strom et al., 2018). Within the ATT&CK framework, the ”Techniques” section comprehensively outlines the specific methods adversaries employ to achieve their tactical objectives. By examining the response rate and error causes for each technique, we gain deeper insights into the performance capabilities and potential limitations of HoneyGPT under different API supports.
The test results, as detailed in Appendix B, indicate that HoneyGPT demonstrates excellent support in responding to each attack technique, achieving response rates above 99%. Notably, the performance of GPT-4 surpasses that of GPT-3.5 Turbo. This analysis not only highlights the robustness of HoneyGPT’s design in simulating realistic and responsive interaction, but also underscores the improvements in handling complex security scenarios facilitated by advancements from GPT-3.5 to GPT-4.
Figure 6. Analyse of GPT-3.5 Turbo Non-Responses
Figure 7. Analyse of GPT-4 Non-Responses
Additionally, Figures 6 and  7 provide an analysis of the causes and frequencies of non-responses encountered by each version of HoneyGPT against various ATT&CK techniques. A notable observation is substantial reductions in commonly encountered error types with GPT-4, such as ”response wrong format”, ”wrong command”, and ”prompt length exceeds limit”, in comparison to GPT-3.5 Turbo.
Response wrong format indicates issues with the parsing of responses due to discrepancies between the returned format and the expected format.
Wrong command occurs when the command is deemed incorrect or inexecutable within the specified system environment.
Prompt length exceeds limit arises when the length of an attack command or response surpasses the LLM’s context length limit, making it impossible to work.
These issues are not easily rectifiable solely through prompt engineering in circumstances where the model’s performance is subpar. By contrast, with the advent of GPT-4, these error types become virtually nonexistent, with only a minuscule number of ”security policy” errors remaining. It highlights the improvement in error management with the introduction of GPT-4, reflecting significant advancements in better interaction than GPT-3.5 Turbo.
Security policy occurs when the ChatGPT refrains from executing a command based on its internal security protocols, and it primarily involves operations such as malicious file downloads or deletions. This type of error is unavoidable in the design of honeypots based on commercial LLMs. Fortunately, at this moment, the frequency of such an error is very low, and their impact on honeypot performance is negligible.
5.4. Baseline Comparison of Flexibility
The evaluation of honeypot flexibility focuses on three key aspects: scalability across different systems, adaptability in different system configurations, and integration of additional security analysis features. The comparative evaluation is among HoneyGPT, Cowrie, Honeyd, and real-system honeypot. Cowrie operates as a medium-interaction honeypot that provides interactive terminal emulation, and Honeyd serves as a low-interaction honeypot designed for general purposes. Real-system honeypots, on the other hand, are actual operating systems running on virtual machines.
Table 5. Flexibility in Simulating Different Systems
Honeypot Expansion Cost Simulation Level Scalability
Cowrie High Medium Medium
Honeyd Low Low High
Real System High High Low
HoneyGPT Low High High
5.4.1. Flexibility in Simulating Different Systems
The diversity of operating systems requires traditional honeypots to adapt to various system versions and instruction sets. This adaptation frequently demands independent development efforts for each simulated system. However, the advanced generative capabilities of LLMs offer a groundbreaking solution to overcome these challenges. We evaluate the emulation capabilities of these honeypots across different operating systems, including Ubuntu and CentOS. Table 5 lists the expansion costs, simulation levels, and scalability of each honeypot type. Cowrie, as a medium-interaction honeypot, requires extensive modifications to adapt to different system versions, including updates to system information and the file systems, as well as reconfigurations of the instruction set, and thus the expansion costs are considerably high. Honeyd, as a low-interaction honeypot, requires only minimal adjustments in configuration information to accommodate system version changes, offering high scalability. Real-system honeypots, while achieving high simulation fidelity, require the configuration of new physical environments, which significantly increases the associated costs and reduces overall flexibility. HoneyGPT requires only minimal adjustments to system settings in prompt to simulate different systems. The LLMs can automatically generate file system, instruction set, system user, and other details based on the system settings in prompt. Our approach offers high scalability and a superior level of simulation.
Table 6. Evaluation of Configuration Flexibility for Different Configuration
Honeypot Type Network Configuration File System User Configuration Hardware Settings Service Status
Cowrie $\times$ $\times$
HoneyD $\times$ $\times$ $\times$
Real-system honeypot $\times$ $\times$ $\times$
HoneyGPT
5.4.2. Flexibility in Simulating Different System Configurations
To assess the flexibility of configuration modifications, we conduct a comparative analysis of Cowrie, Honeyd, real-system Honeypot, and HoneyGPT. We categorize configuration changes into five types: network configuration, file system, user configuration, hardware setting, and service status. If configuration updates can be easily achieved by simple command-line or configuration file modifications, we consider it as supporting flexible modification for that configuration. Table 6 shows that HoneyGPT is the only product allowing flexible modifications across all configuration aspects. Honeyd, Cowrie, and real-system honeypots exhibit limited configuration flexibility.
Traditional honeypots often struggle with modification of system configurations, leading to fixed deployments that can be easily identified by attackers. This rigidity undermines their effectiveness as deception tools. However, the advent of LLMs introduces a novel paradigm. HoneyGPT enables direct modifications to system configurations via prompt adjustments, offering unprecedented flexibility and adaptability, thereby enhancing deception capabilities.
5.4.3. Flexibility of Integration with Security Analysis Capability
In the context of practical operations, honeypots primarily act as collectors of attack data, representing just one facet of comprehensive security measures. Beyond this basic role, the analysis of collected attack data is crucial. Traditional honeypot designs, such as Cowrie, Honeyd, require specialized development efforts, thereby incurring significant overheads to integrate these functionalities.
However, the HoneyGPT’s design circumvents this challenge by leveraging the capabilities of ChatGPT to seamlessly integrate additional functionalities through simple prompts. For example, we can employ ChatGPT to analyze attackers’ behavioral logs, enabling functionalities like security policy alerts. As designed in the prompt manager of HoneyGPT, each interaction with LLMs not only inquires about the outcome of the current command but also asks about the potential impact of the attack command on the operating system.
This innovative approach represents a paradigm shift where a honeypot, by leveraging LLMs, seamlessly integrates security analysis capabilities. The utility of the honeypot is greatly enhanced while reducing the cost and complexity associated with the integration of these auxiliary features. Therefore, HoneyGPT presents a cost-effective, scalable, and efficient approach for future honeypots.
5.5. Field Evaluation
The field evaluation focuses on assessing HoneyGPT’s entrapment efficacy in the real world. We have simultaneously deployed HoneyGPT based on GPT-4 and Cowrie inside the same network for four weeks. HoneyGPT is implemented using the best practice deployment strategies, while Cowrie is configured with its default settings. We detail the evaluation procedure and our main findings below.
5.5.1. HoneyGPT Deployment
Our baseline comparisons demonstrate that HoneyGPT significantly enhances deception, interactivity, and flexibility. However, the economic burden associated with the query costs of commercial LLMs presents practical challenges for deployment. Furthermore, these models also impose restrictions on the quantity and frequency of requests from a single user within a given timeframe, potentially limiting their capacity to handle tasks with high concurrent requests. To optimize the cost-effectiveness of HoneyGPT under the constraints of high query costs and stringent request limitations, we adopt a hybrid deployment strategy. This strategy integrates HoneyGPT with the traditional emulated honeypot approach, leveraging the strengths of both to enhance performance while reducing costs.
In the strategic deployment of HoneyGPT, prudent management of command processing is crucial for optimizing efficacy and cost-efficiency. Those commands that are challenging for LLMs to process, as well as those governed by fixed, simple rules, are delegated to emulated honeypots. In other words, the commands that LLMs typically handle inefficiently, such as binary file reading, and those dependent on time mechanisms (e.g., sleep and uptime), as well as commands following simple rules, such as whoami, are rerouted to emulated honeypots.
Furthermore, resolved interactions from HoneyGPT—comprising terminal prompts and command combinations—are stored in a database. If an attacker’s request corresponds to a stored HoneyGPT query, the recorded response from the database is immediately returned. For new, context-sensitive queries, processing is handled by HoneyGPT to ensure responses are coherent and contextually accurate. Due to the repetitiveness of attacker requests in actual intrusion scenarios, this strategy significantly reduces the query load and congestion on commercial LLMs.
Moreover, to enhance the appeal of the honeypot to specific adversaries, such as those involved in cryptocurrency mining, the simulated environment within HoneyGPT is configured to mimic a high-performance computing server. Prompts are carefully crafted to align with attackers’ intentions, allowing for the successful execution of attack commands. This is intended to modify attacker behaviors and elicit more prolonged and informative interactions, thereby providing deeper insight into their strategies and potentially uncovering more complex attack techniques.
5.5.2. Comparative Analysis of Attacker Engagement
By comparing the interaction data of HoneyGPT and Cowrie, we observe that for the same attack source under the three cases, HoneyGPT achieves a deeper interaction with the attacker than Cowrie.
  • Fulfillment of Attacker’s Intent. As illustrated in Figure 8, attackers issue the ps and grep command combination to identify processes related to the ”miner” keyword, aiming to detect mining malware. Cowrie cannot reveal the processes with the ”miner” keyword, leading to a premature termination of the attack. Conversely, HoneyGPT effectively mimics the expected system responses, thus fulfilling the attacker’s intent and encouraging sustained engagement.
    Figure 8. Fulfillment of Attacker’s Intent
  • Command Support Level. As illustrated in Figure 9, when attackers try to use complex command structures with multiple command combinations, the limited interaction capabilities of the Cowrie honeypot often fail to support all the sub-commands within these complex commands. As a result, attackers typically cease their attacks after unsuccessful attempts using similar commands. By contrast, HoneyGPT’s generative capabilities are sufficient to handle such complex commands, which can entice attackers to engage in deeper levels of attacks.
    Figure 9. Command Support Level
  • Content Rigidity. As illustrated in Figure 10, when attackers engage in reconnaissance of the target system, Cowrie typically provides rigid, easily identifiable honeypot responses. Recognizing these as features of a honeypot, attackers often terminate their interaction prematurely. By contrast, HoneyGPT dynamically generates responses tailored to the honeypot setting outlined in the prompt, effectively camouflaging its honeypot nature and maintaining attacker engagement.
Figure 10. Content Rigidity
This comparative analysis highlights HoneyGPT’s superior capabilities in providing dynamic, contextually relevant content that can significantly sway attacker behaviors, support a broader array of command interactions, and effectively disguise its honeypot identity, thereby extending engagement and potentially exposing more advanced attack strategies.
5.5.3. New Attacks Vector Discovered by HoneyGPT
The deployment of HoneyGPT has unveiled new attack vectors that significantly differ from those captured by the traditional Cowrie honeypot. As detailed in Appendix C, HoneyGPT identifies 11 types of attack actions categorized into six different ATT&CK techniques, highlighting the evolving landscape of cyber threats and the critical role of advanced, adaptive honeypot technologies like HoneyGPT in identifying and analyzing emerging attack vectors.
6. Limitations and Future Work
While the integration of LLMs has enhanced the entrapment capabilities of honeypot systems, there remain several limitations manifested in the following aspects:
  • Context Length Restriction. LLMs are inherently limited by the Context length of prompt. When attackers input and execute lengthy command scripts directly at the terminal, LLMs may fail to respond appropriately. Attempts to limit response lengths within the prompts have proven suboptimal. This issue may necessitate the development of proprietary LLMs tailored for extensive token interactions.
  • Request Rate Limit. Although the use of rule-based matching has reduced the amount of unnecessary commercial LLM requests, commercial models’ limitations on request rates remain a significant obstacle to the deployment of honeypot systems. Professional honeypot models, fine-tuned from open-source LLMs, could potentially address this issue.
  • Ineffectiveness against Fixed Attack Sequences. During field testing, we discovered that some attackers employ predetermined attack sequences to target honeypots, regardless of the honeypot’s ability to respond to requests or the quality of the response content. For such attackers, HoneyGPT is unable to prolong the interaction length by catering to their intent.
  • Prompt-based Attacks on ChatGPT. While no prompt-based attacks against ChatGPT have been discovered thus far, this remains a potential vulnerability that needs to be monitored and addressed (Si et al., 2023; Greshake et al., 2023).
  • Overfitting Issue. LLMs may exhibit overfitting to incorrect commands, returning seemingly normal results despite command errors. This issue highlights the need for further refinement in the model’s error recognition capabilities.
Moving forward, the development of LLMs for honeypot applications should focus on mitigating these limitations. Research efforts may include creating more sophisticated and context-aware models that can handle longer token sequences and have a deeper understanding of complex systems. Furthermore, enhancing models with temporal awareness and addressing request rate limits will be crucial for advancing honeypot technology. It will also be paramount to continually monitor for new types of attacks and adapt the LLMs accordingly to maintain robust security measures. Finally, addressing the overfitting problem may involve implementing more nuanced training strategies that emphasize the correct handling of anomalous or incorrect command inputs.
7. Conclusion
The triad of challenges in traditional honeypots—flexibility, interaction depth, and deceptive capacity—has historically curtailed their allure and efficacy in engaging attackers. To address this trilemma, this paper presents HoneyGPT, a bespoke innovation in adaptive honeypot solutions specifically engineered to circumvent the inherent limitations of traditional honeypots. Our comparison evaluations indicate that HoneyGPT transcends traditional honeypots in terms of flexibility, interaction, and deception, and it also exceeds the entrapment capabilities of real systems, emphasizing its superiority in meeting attackers’ intents to deepen engagements. Our field evaluations further validate that HoneyGPT adeptly conducts extended interactions and captures a broader spectrum of attack vectors while facing diverse adversaries in the real world. Overall, HoneyGPT positions itself as an effective and intelligent countermeasure against ongoing security threats, and it also paves the path for future honeypot research.
Appendix A Configuration of HoneyGPT
A.1. System Principles $P$
HoneyGPT, serving as a terminal honeypot system, is specifically designed to attract and monitor attack activities. To achieve this objective, we have implemented the system’s principles within the prompts, providing operational guidelines for ChatGPT. These prompts play different crucial roles in the functionality of the system, including:
Role of HoneyGPT
HoneyGPT strategically simulates an authentic terminal environment, tailored to align closely with attackers’ intents, thereby encouraging deeper and more sustained engagement. The system expertly balances authenticity with the objective of enticing attackers into prolonged interactions, thereby enhancing the honeypot’s effectiveness.
Time Sensitivity
In scenarios where attacks are time-sensitive, attackers may issue commands related to current time inquiries, such as uptime and top. Since ChatGPT inherently lacks the capability to query real-time networked time data, it is necessary for us to provide time-related information such as the current time and boot time in the prompts to ChatGPT. This approach ensures that the honeypot’s responses are temporally accurate, thus more convincingly mimicking a real system environment.
Input/Output Format
The structure of both input and output prompts critically defines each interaction with the LLM in HoneyGPT. The input format is crucial for the LLM’s accurate understanding of the honeypot system’s state and the commands issued by attackers. Conversely, the output format affects HoneyGPT’s operational effectiveness. To ensure HoneyGPT’s robustness, we instruct the LLM to use the JSON format for outputs. This standardization ensures consistent and correct parsing of responses during each interaction with the model.
Few-Shot Learning
HoneyGPT utilizes a few-shot learning strategy to enhance ChatGPT’s comprehension of task requirements and improve the quality of its outputs, compensating for gaps in several command knowledge. Our experience indicates that by integrating only 4-5 representative examples, HoneyGPT’s generative capabilities can be significantly enhanced.
A.2. Honeypot Setting $S$
Honeypot settings describe the terminal system information simulated by HoneyGPT, which is crucial for attracting and capturing potential attackers. Our experience indicates that honeypots with high-end configurations are more attractive to malicious entities, such as cryptocurrency-mining malware. Furthermore, the more details of honeypot configurations, the higher precision of the responses. The ability to customize honeypot settings via natural language significantly enhances flexibility and reduces the complexity of deployment. Honeypot settings are divided into two main categories: hardware and software.
Hardware
It is essential to define the hardware specifications that will be emulated by HoneyGPT. It includes defining CPU types and counts, GPU presence, and storage capacities, those attributes that attackers frequently scrutinize. The specifications selected must mirror those of the systems intended to be simulated by HoneyGPT to maintain the authenticity of the honeypot environment. Additionally, configuring systems with high-end GPUs and CPUs can attract specific types of threat actors, such as crypto-mining malware, prompting deeper interactions.
Software
The software environment of a honeypot, including the operating system, open ports, user configurations, and application services, must be strategically crafted to reflect the behaviors of potential targets. For instance, if the objective is to mimic a web server, HoneyGPT should be configured with the appropriate web server software, along with associated services and open ports. Additionally, attackers are often drawn to details such as system processes, resource utilization, scheduled tasks, user configurations, and filesystem information. These elements can be specifically tailored within HoneyGPT to create customized responses for each engagement scenario.
Appendix B Successful Response Rate of HoneyGPT based on GPT-3.5-turbo and GPT-4
Figure 11 shows the successful response rate of HoneyGPT based on GPT-3.5-turbo and GPT-4, with support rates for each attack technique exceeding 99%.
Figure 11. Successful Response Rate with GPT-3.5 Turbo and GPT-4 in HoneyGPT
Appendix C New Attacks Vector Discovered by HoneyGPT
As shown in Table 7, HoneyGPT identifies 11 types of attack actions, which are classified into 6 categories of ATT&CK techniques.
Table 7. New Attacks Vector Discovered by HoneyGPT
Technology action
System Service Discovery (T1007) By employing pipeline operators (—) to integrate the nvidia-smi command with grep and other filtering commands like head, awk, and wc, these commands facilitate the selective extraction and presentation of precise GPU information in accordance with predetermined criteria.
System Service Discovery (T1007) By employing pipeline operators (—) to merge the lspci command with various filtering commands including grep, cut, egrep, and wc. The objective is to extract and display specific information related to VGA or 3D graphics from the PCI devices present within the system.
System Service Discovery (T1007) The lscpu command is integrated with egrep and cut to selectively extract the CPU model name.
System Service Discovery (T1007) The grep command is utilized to search the /proc/cpuinfo file to calculate the total number of processors or CPU cores in the system.
System Network Configuration Discovery (T1016) The command arp -a is used to display the system’s ARP table, containing mappings between IP addresses and MAC addresses.
System Network Configuration Discovery (T1016) An HTTP GET request is sent using curl to ”curl ipinfo.io/org” in order to retrieve organizational information associated with the system’s IP address that initiated the request.
System Network Configuration Discovery (T1016) The nano command is used to view the /etc/resolv.conf file, to access the configuration of the DNS resolver.
System Information Discovery (T1082) The uptime command is employed to obtain the system’s uptime.
Indicator Removal on Host (T1070) Several actions are taken to ensure anonymity and to avoid leaving traces of system activity: 1. Unsetting environmental variables to disable the saving of command history. 2. The command history -c is used to clear the existing history. 3. The rm -rf command combined with the path to log files is used to delete log files, thereby eliminating traces of system activity.
Virtualization/Sandbox Evasion( T1497) The sleep command is used to nsuring program synchronization, preventing detection of anomalous behavior, and it can also be used to detect if the target system is a honeypot, in which case, if execution fails, further attacks are not pursued.
Obfuscated Files or Information (T1027) Scripts written in shell or Perl are obfuscated by base64 encryption prior to execution to avoid detection and identification.